OpenStudy (anonymous):

-2x - 18 = 6(1 - 2x) + 4x x =

OpenStudy (anonymous):

first, multiply out (or distribute) the 6(1-2x) on the right side.

OpenStudy (anonymous):

so..... -2x-18=6-12x+4x

OpenStudy (anonymous):

combine the like terms -2x-18=6-8x

OpenStudy (anonymous):

then do -2x-18=6-8x +2x +2x -18=6-6x -6 -6 -24/6=-6x/6 4=x

OpenStudy (anonymous):

medal?? :)